Project

General

Profile

Kaon LT Meetings » mtg_22sep21.txt

Garth Huber, 09/21/2022 05:50 PM

 
1
                Sept 21/22 PionLT/KaonLT Analysis Meeting Notes
2
                -----------------------------------------------
3
                           (Notes by GH and SJDK)
4

    
5
                    Today: KaonLT will be discussed first
6

    
7
Please remember to post your slides at:
8
https://redmine.jlab.org/projects/kltexp/wiki/Kaon_LT_Meetings
9

    
10
Present:
11
Regina - Stephen Kay, Garth Huber, Nathan Heinrich, Vijay Kumar, Alicia Postuma
12
   Muhammad Junaid, Love Preet, Ali Usman
13
FIU - Pete Markowitz
14
Ohio - Jacob Murphy, Julie Roche
15
CSULA - Konrad Aniol, Emery Mark Gunselman, Erika Gwin
16
CUA - Richard Trotta, Tanja Horn
17
Glasgow - David Hamilton
18

    
19
Discussion before we start:
20
   - CaFe is running a different trigger setup. Coin and 3/4 arriving at the
21
     same time?
22
   - GH is interested in finding out about the data they took for "proton
23
     absorption correction" and "Heep normalization"
24

    
25
- Konrad introduces Emery and Erika, two 1st year Masters students who've just
26
  started.  Cal state has an interest in Hall A as well as Hall C
27

    
28
Richard Trotta Updates
29
----------------------
30
- KaonLT Lumi Analysis
31
   - Scaler yield calculation
32
      - using pTrig3=ELREAL
33
      - Had CPULT applied.  Now removed and things look better, although 
34
        scaler yield still increases with current
35
   - Cuts shown for Tracked vs Non-Tracked analyses
36
      - Cuts are a little on the tight side
37
      - Cuts are all lower bound, no upper bound cuts, so should be okay
38
   - Tracked yield calculation has the total efficiency
39
      - Tracking efficiencies in report file
40
   - Jacob points out that we could have Evttype == 3 too
41
   - Tracked yield drops very quickly with current, definitely an issue!
42
      - How old is the replay?  From May
43
   - *Reminder: *SHMS was pi+ polarity for one scan, e- for everything else so
44
       the PID cuts used for the pi+ Lumi scan will be much different than for
45
       all other scans
46
      - -ve polarity lumi scan runs have low TLT
47
      - +ve polarity lumi scans have some issues with the cuts included in the
48
        replay
49
         - planning to implement Aerogel Tray cut for +ve runs
50
   - Seem to be a couple of issues tied to replay
51
      - 6p2 GeV Lumi scans - very low TLT
52
      - Low beam on time.  Possibly tied to the >5uA current cut?
53
      - Could also be linked to cuts defined in the DEF-CUTS file
54
   - GH: Please add a plot of the actual DAQ rate so we can see whether low
55
     CPULT and TLT makes sense or not
56
   - Ali: made subtle changes to the event selection to enable tracking
57
     efficiencies to be calculated for multiple event types
58
      - was restricting event type after decode EDTM info
59
      - now doing ALL events
60
      - not sure whether this will affect Lumi analysis or not
61

    
62
- Bill's LT-sepaaration section code
63
   - Richard going through code to understand and get it working
64
      - Original code was a little messy, unclear variables etc
65
      - Lots of hard coding, particularly in analysis.cpp
66
      - Setting up pathing for filenames after analysis
67
   - it is inevitable that some hard coding will remain, since the number of
68
     SHMS angles is different at high/low epsilon for different settings
69
      - making backend scripts for plotting that are separate for each setting
70
        that can be called by master scripts as necessary
71
   - GH/TH ask him to post some notes as part of code documentation, kaonlt log
72
      - RT will post notes in sections as they are completed
73
   - will try to have a working version to show in 2wks
74
   - GH will go again through his LT-separation tutorial next Tues @ 12:00 EDT
75
   
76
- To do/upcoming
77
   - Calorimeter calibrations
78
   - HGCer efficiency calculation
79
      - Ali has some notes on this
80

    
81
Ali Usman Updates
82
-----------------
83
- Total efficiency per run
84
   - All efficiencies and LTs multiplied together
85
      - HMS ElTrack
86
      - SHMS ProtonTrack
87
      - Hodo 3/4 (both)
88
         - Need to verify how hcana actually calculates the hodoscope
89
	   efficiencies
90
      - EDTM LT
91
      - HMS Cherenkov
92
   - Effective Q=Total Eff * Q
93
      - sum Effective Q for all runs in a setting
94
      - yield Y=Counts/Effective Q
95

    
96
- Heep Coin Data/SIMC comparisons
97
   - with latest replay, gbeam.parm fixed
98
   - Comparison looks good, but plots include too many types of data
99
     (unsubtracted, w/o cuts, etc) and this partly obscures the comparison
100
      - Future versions should only be the two things we actually want to
101
        compare
102
   - Data 3% low compared to MC
103
      - still need to apply cryotarget boiling and proton absorption, so we
104
        might eventually with data a bit too high compared to MC
105
   - Need to finalise offsets for KaonLT data (all of it)
106
      - Ali and Vijay will work together to try to find offsets that aren't too
107
        different from each other, if possible
108

    
109
- Tanja shares the Fpi-2 logbook
110
   - https://hallcweb.jlab.org/fpilog/
111
     Username: hms  Password: sos
112

    
113
- Fpi-2 website
114
   - https://hallcweb.jlab.org/experiments/fpi/FPI_PHASE2/
115

    
116
Vijay Kumar Updates
117
-------------------
118
- Updated gbeam.param file
119
   - Switched to one Dave G had recommended
120
   - Should now have consistent param files with Ali/Richard
121

    
122
- Working on Lumi study
123
   - Low energy PionLT data
124
   - Ran into some issues with 
125

    
126
- Summer 2019 data gbeam.param?
127
   - Do we need a new file?
128
   - Dave G has a script that gives some of the beam on target parameters that
129
     go in the file
130
   - it should be the same as KaonLT except for this change
131

    
132
Stephen Kay Updates
133
-------------------
134
- Merged in new online updates to offline branch
135
   - Should not be any conflicts
136
      - Scripts are compartmentalised in UTIL directories
137
      - Merger doesn't cause any issues with this
138

    
139
- Should start to organise PionLT param and DB files as we did with KaonLT
140
   - Ali did some organization of KaonLT param
141
      - every detector is organized in directories
142
      - better isolates for chagnes, to avoid overwriting
143
   - also need to trim down std.kin, just PionLT settings
144
   - Nathan: we should make these changes as we begin doing the calibs
145

    
146
Jacob Murphy Updates
147
--------------------
148
- will start with HGC, NGC calibs
149

    
150
- HMS Optics work
151
   - first optics pass for HMS 6.8GeV calibrations
152
   - Small issue where it's not getting the correct ytar value?
153
   - Look at where holes end up from all three foils
154
   - Jacob demonstrates why we meed the 6 GeV optical matrix
155
      - Peaks in H.gtr.y blend into each other if the wrong one is used
156
      - Ysieve resolution much better than Xsieve for some settings
157
      - possibly saturation effects larger for Ysieve?
158
      
159
Next meeting: Wed Sept 28 @ 9:00 Eastern/7:00 Regina/6:00 Calif
160
   - Jacob will start first next time
161
   - ask to give some intro on the optics fitting for those less familiar
(92-92/430)